# Robust targeted exploration for systems with non-stochastic disturbances

### v2

@@ 5 Numerical Example
Comparison with naive exploration: To highlight the benefits of the proposed method, we compare the targeted exploration method with a naive exploration strategy that uses inputs of equal total energy but with non-optimized amplitudes, i.e., inputs where the energy is uniformly distributed across all frequencies. This baseline does not rely on any model information and illustrates how optimization improves the input design. We select the same eight uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}},5\times 10^{i}I_{n_{\phi}}\}$, i=4,,7i=4,...,7 as above, and $\gamma_{\mathrm{w}}=1$. For each trial, we augment the datasets with data generated using the corresponding naive input. From the datasets, we compute the estimate of the parameters $\hat{\theta}_{T}$, and $\|G\cdot P\|$ for the proposed targeted exploration and the baseline. From Figure 3, it can be observed that the proposed targeted exploration method guarantees a lower error bound $\|G\cdot P\|$, by roughly $50\%$, compared to naive exploration. This comparison demonstrates that the proposed targeted input design guarantees lower parameter error under the same energy budget, thereby highlighting the benefits of our approach.
@@
Figure 3: Illustration of the a posteriori guaranteed bound on the squared error of the parameters $\|G\cdot P\|$ for both targeted and naive exploration with same input energy, for different initial uncertainty bounds $D_{0}^{-1}$.
@@
Sensitivity related to the initial estimate $\hat{\theta}_{0}$: To study the sensitivity related to the initial estimate $\hat{\theta}_{0}$, we select three different uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}}\}$, $i=4,5,6$. We select $\gamma_{\mathrm{w}}=1$. For each uncertainty level, we run 50 trials. Each trial comprises: $\mathrm{(i)}$ generating a random initial estimate $\hat{\theta}_{0}$ that satisfies Assumption 2, i.e., $(\hat{\theta}_{0}-\theta_{\mathrm{tr}})^{\top}(D_{0}\otimes I_{n_{\mathrm{x}}})(\hat{\theta}_{0}-\theta_{\mathrm{tr}})\leq 1$, $\mathrm{(ii)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$, $\mathrm{(iii)}$ executing Algorithm 1 to obtain the exploration inputs (6) and the input energy $\gamma_{\mathrm{e}}^{2}$.
@@
From Figure 4, it can be observed that the variability of input energy $\gamma_{\mathrm{e}}^{2}$ is higher for larger initial uncertainty levels $\|D_{0}^{-1}\|$. This variability of the input energy significantly reduces as initial uncertainty levels reduce. This is because small initial uncertainty levels yield a tighter range of bounds on $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, thereby leading to more consistent exploration input energy.
Sensitivity related to the desired error bound $\|D_{\mathrm{des}}\|$: To study the sensitivity related to the desired uncertainty level $\|D_{\mathrm{des}}\|$, we select six levels of uncertainty $D_{\mathrm{des}}=10^{i}I_{n_{\phi}}$, $i\in\{-4,-3,-2,-1,0,1\}$. We set the initial uncertainty level to $D_{0}=10^{4}I_{n_{\phi}}$, $\gamma_{\mathrm{w}}=1$ and select the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$. We compute the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$. For each desired uncertainty level $\|D_{\mathrm{des}}\|$, we execute Algorithm 1 to obtain the exploration inputs (6) and the input energy $\gamma_{\mathrm{e}}^{2}$.
@@
From Figure 5, it can be observed that achieving a small uncertainty (large $\|D_{\mathrm{des}}\|$) requires larger input energy. In contrast, in the setting of stochastic disturbances [venkatasubramanian2023sequential], the required input energy scales linearly with $\|D_{\mathrm{des}}\|$. This difference is due to the data-dependent scaling $G$ (14) of the non-falsified set for energy-bounded disturbances (cf. Lemma 6).
Computational time: The average execution time for solving Problem (4.6) over 10 trials is approximately 45 seconds. In general, Problem (4.6) is a standard SDP, and its computational effort scales polynomially with the state dimension $n_{\mathrm{x}}$, control input dimension $n_{\mathrm{u}}$, and the number of frequencies $L$. While this makes moderately-sized problems tractable, the computational effort may become prohibitive for large-scale problems. However, for large-scale problems, problem-specific structures can often be exploited and efficient solvers may be utilized to reduce computational demand [majumdar2020recent]. The simulations were carried out on a system with an AMD Ryzen 7 5700U processor and 16.0 GB RAM.
In summary, the simulation results demonstrate the applicability of the proposed exploration strategy to systems with unmodeled nonlinearities. Given an initial estimate $\hat{\theta}_{0}$ and initial uncertainty bound $D_{0}^{-1}$, the targeted exploration strategy guarantees an a priori error bound on parameters estimated after exploration.
@@ 6 Conclusion
In this article, we presented a targeted exploration strategy for linear systems subject to energy-bounded disturbances. We derived LMIs that can robustly guarantee an a priori error-bound on the estimated parameters after exploration. The proposed strategy utilizes multi-sine inputs in selected frequencies and optimized amplitudes to shape the uncertainty bound over the parameters in a targeted manner. We have demonstrated the applicability of the proposed targeted exploration strategy to systems with nonlinearities, while analysing its benefits and inherent conservatism, with a numerical example. Overall, to the best of the authors’ knowledge, this is the first targeted exploration approach that robustly ensures a user-chosen accuracy on the parameters without requiring any independence conditions on the disturbances.
@@ Appendix A Proof of Lemma 9
By Young’s inequality [caverly2019lmi], for any $\epsilon>0$, we have
@@
By Young’s inequality [caverly2019lmi], for any $\epsilon>0$, we have
@@ Appendix B Proof of Lemma 10
Appendix C Proof of Theorem 14
@@
Inequalities (48a)-(48c) imply the exploration goal (4) due to Proposition 13. In what follows, we prove that Inequalities (49a)-(49c) imply Inequalities (48a)-(48c), respectively. In particular, we utilize the matrix S-lemma [boyd2004convex, vanwaarde2022noisy] to account for uncertainties in $V_{\mathrm{x,tr}}$, $V_{\mathrm{\phi,tr}}$, $Y_{\mathrm{x,tr}}$, and $Y_{\mathrm{\phi,tr}}$ satisfying bounds (38), (39). The proof is divided into three parts wherein each part derives (49a), (49b) and (49c), respectively.
@@
By using the matrix S-lemma [boyd2004convex, vanwaarde2022noisy], Inequality (C) holds for all $V_{\mathrm{x,tr}}$ satisfying Inequality (C), if $S_{\textnormal{exp-1}}(\epsilon,\tau_{1},U_{\mathrm{e}},\hat{Z},\hat{V}_{\mathrm{x}},\tilde{\Gamma}_{\mathrm{x}},\tilde{D}_{\mathrm{des}},\bar{D}_{1})~\succeq~0$ (49a) holds with $\tau_{1}\geq 0$.
@@
By using the matrix S-lemma, Inequality (C) holds for all $V_{\mathrm{\phi,tr}}$ satisfying Inequality (C), if $S_{\textnormal{exp-2}}(\epsilon,\tau_{2},U_{\mathrm{e}},\hat{Z},\bar{W}_{Z},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},\tilde{D}_{\mathrm{des}},\bar{D}_{2})\succeq 0$ (49b) holds with $\tau_{2}\geq 0$.
@@
By using the matrix S-lemma, Inequality (C) holds for all $V_{\mathrm{\phi,tr}}$ satisfying Inequality (C), if the following holds with $\tau_{3}\geq 0$:
@@
From Lemma 12, we have
@@
Inserting Inequality (80) in Inequality (C) yields $S_{\textnormal{exp-3}}(\epsilon,\tau_{3},U_{\mathrm{e}},\hat{U},\hat{W}_{\phi},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},D_{\mathrm{des}},\bar{D}_{3})\succeq 0$ (49c).

### v3

@@ 5 Numerical Example
Conservatism related to the initial uncertainty bound $D_{0}$ and comparison with naive exploration: To study the effect of the initial uncertainty bound $D_{0}^{-1}$, we select eight uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}},5\times 10^{i}I_{n_{\phi}}\}$, i=4,,7i=4,...,7 and $\gamma_{\mathrm{w}}=1$. Based on $\gamma_{\mathrm{w}}$, we compute $\zeta_{1}=\zeta_{2}=\frac{m_{1}m_{2}\sqrt{\gamma_{\mathrm{w}}}}{\sqrt{T(m_{1}^{2}+m_{2}^{2})}}$. Furthermore, to highlight the benefits of the proposed method, we compare the targeted exploration method with a naive exploration strategy that uses inputs of equal total energy but with non-optimized amplitudes, i.e., inputs where the energy is uniformly distributed across all frequencies. This baseline does not rely on any model information and illustrates how optimization improves input design. Each trial comprises: $\mathrm{(i)}$ computing the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$, $\mathrm{(ii)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, $\mathrm{(iii)}$ executing Algorithm 1 to obtain the exploration inputs (6), and generate a dataset by applying the targeted exploration input, $\mathrm{(iv)}$ computing the naive exploration inputs, and generating a corresponding dataset. From both datasets, we compute the corresponding estimate of the parameters $\hat{\theta}_{T}$, and $\|G\cdot P\|$, the error bound guaranteed by Lemma 6 with the data-dependent covariance matrix $P$ (9) and scaling $G$ (11).
@@
Figure 2: Illustration of the a posteriori guaranteed bound on the squared error of the parameters $\|G\cdot P\|$ for both targeted and naive exploration with same input energy, in comparison with the desired bound on the squared error $\|D_{\mathrm{des}}^{-1}\|$ for different initial uncertainty bounds $D_{0}^{-1}$.
Figure 2 shows that the targeted exploration inputs achieve the exploration goal for all tested initial uncertainty levels $D_{0}$ and energy bound $\gamma_{\mathrm{w}}$, i.e., the guaranteed error bound on the parameters $\|G\cdot P\|$ is lower than the desired error bound $\|D_{\mathrm{des}}^{-1}\|~=~10^{0}$. The guaranteed bound is significantly lower than the desired bound due to the inherent conservatism of the proposed strategy which is derived by utilizing worst case bounds on the transfer matrices $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, and worst case bounds $\bar{W}_{\phi}$, $\bar{W}_{Z}$. This conservatism decreases as the initial uncertainty is reduced, or equivalently, as $D_{0}$ increases, since smaller initial uncertainty results in tighter and more accurate bounds on $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, yielding more efficient exploration. Furthermore, Figure 2 shows that the proposed targeted exploration method guarantees a lower error bound $\|G\cdot P\|$, by roughly $50\%$, compared to naive exploration under the same energy budget, thereby highlighting the benefits of our approach.
@@
Sensitivity related to the initial estimate $\hat{\theta}_{0}$: To study the sensitivity related to the initial estimate $\hat{\theta}_{0}$, we select three different uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}}\}$, $i=4,5,6$, with $\gamma_{\mathrm{w}}=1$. For each level, we run 50 trials. Each trial comprises: $\mathrm{(i)}$ generating a random initial estimate $\hat{\theta}_{0}$ that satisfies Assumption 2, i.e., $(\hat{\theta}_{0}-\theta_{\mathrm{tr}})^{\top}(D_{0}\otimes I_{n_{\mathrm{x}}})(\hat{\theta}_{0}-\theta_{\mathrm{tr}})\leq 1$, $\mathrm{(ii)}$ computing constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, $\mathrm{(iii)}$ executing Algorithm 1 to obtain the exploration inputs (6) and the input energy $\gamma_{\mathrm{e}}^{2}$.
@@
Figure 3 shows that the variability of input energy $\gamma_{\mathrm{e}}^{2}$ is higher for larger initial uncertainty levels $\|D_{0}^{-1}\|$. This variability significantly reduces as initial uncertainty levels reduce. This behaviour arises because smaller initial uncertainty levels yield tighter bounds on $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, resulting in more consistent input energy.
Sensitivity related to the desired error bound $\|D_{\mathrm{des}}\|$: To study the sensitivity related to the desired uncertainty level $\|D_{\mathrm{des}}\|$, we select six levels $D_{\mathrm{des}}=10^{i}I_{n_{\phi}}$, $i\in\{-4,-3,-2,-1,0,1\}$. We set the initial uncertainty level to $D_{0}=10^{4}I_{n_{\phi}}$, $\gamma_{\mathrm{w}}=1$, choose $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$, and compute the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$. For each $\|D_{\mathrm{des}}\|$, we execute Algorithm 1 to obtain the exploration inputs (6) and the input energy $\gamma_{\mathrm{e}}^{2}$.
@@
Figure 4 shows that achieving a small uncertainty (large $\|D_{\mathrm{des}}\|$) requires higher input energy. In contrast, the required input energy scales linearly with $\|D_{\mathrm{des}}\|$ in the stochastic disturbances setting [26]. This difference is due to the data-dependent scaling $G$ (11) of the non-falsified set $\mathbf{\Theta}_{T}$ (cf. Lemma 6).
Computational time: The average execution time for solving Problem (4.6) over 10 trials is approximately 45 seconds. The simulations were carried out on a system with an AMD Ryzen 7 5700U processor and 16.0 GB RAM.
@@ 6 Conclusion
In this article, we present a targeted exploration strategy for linear systems subject to energy-bounded disturbances. We derived LMIs that can robustly guarantee an a priori error-bound on the estimated parameters after exploration. The proposed strategy utilizes multi-sine inputs in selected frequencies and optimized amplitudes to shape the uncertainty bound over the parameters in a targeted manner. We demonstrate the applicability of the proposed targeted exploration strategy to systems with nonlinearities, while analysing its benefits and inherent conservatism, with a numerical example. This is the first targeted exploration approach that robustly ensures a user-chosen accuracy on the parameters without requiring any independence conditions on the disturbances. We view addressing the limitations related to conservatism and scalability as an important direction for future work.
@@ Appendix A Proof of Lemma 9
By Young’s inequality [6], for any $\epsilon>0$, we have
@@
By Young’s inequality [6], for any $\epsilon>0$, we have
@@ Appendix B Proof of Lemma 10
Appendix C Proof of Proposition 11
@@
Starting from Inequality (11), we have
[…]
Appendix D Proof of Proposition 12
{pf}
For matrices $\bar{Z}_{\mathrm{u}}$ and $\hat{Z}$, we have $\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\bar{Z}_{\mathrm{u}}\hat{Z}^{\mathsf{H}}-\hat{Z}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}+\hat{Z}\hat{Z}^{\mathsf{H}}=(\bar{Z}_{\mathrm{u}}-\hat{Z})(\bar{Z}_{\mathrm{u}}-\hat{Z})^{\mathsf{H}}\succeq 0$, and hence, we have
[…]
Appendix E Proof of Theorem 13
{pf}
Inequalities (42a)-(42c) imply the exploration goal (4) due to Proposition 12. In what follows, we prove that Inequalities (43a)-(43c) imply Inequalities (42a)-(42c), respectively. In particular, we utilize the matrix S-lemma [5, 24] to account for uncertainties in $V_{\mathrm{x,tr}}$, $V_{\mathrm{\phi,tr}}$, $Y_{\mathrm{x,tr}}$, and $Y_{\mathrm{\phi,tr}}$ satisfying bounds (35), (36). The proof is divided into three parts wherein each part derives (43a), (43b) and (43c), respectively.
@@ Appendix E Proof of Theorem 13
By using the matrix S-lemma [5, 24], Inequality (E) holds for all $V_{\mathrm{x,tr}}$ satisfying Inequality (E), if $S_{\textnormal{exp-1}}(\epsilon,\tau_{1},U_{\mathrm{e}},\hat{Z},\hat{V}_{\mathrm{x}},\tilde{\Gamma}_{\mathrm{x}},\tilde{D}_{\mathrm{des}},\bar{D}_{1})~\succeq~0$ (43a) holds with $\tau_{1}\geq 0$.
@@
By using the matrix S-lemma, Inequality (E) holds for all $V_{\mathrm{\phi,tr}}$ satisfying Inequality (E), if $S_{\textnormal{exp-2}}(\epsilon,\tau_{2},U_{\mathrm{e}},\hat{Z},\bar{W}_{Z},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},\tilde{D}_{\mathrm{des}},\bar{D}_{2})\succeq 0$ (43b) holds with $\tau_{2}\geq 0$.
@@
By using the matrix S-lemma, Inequality (E) holds for all $V_{\mathrm{\phi,tr}}$ satisfying Inequality (E), if the following holds with $\tau_{3}\geq 0$:
@@
For matrices $U_{\mathrm{e}}$ and $\hat{U}$, we have $\bar{U}_{\mathrm{e}}\bar{U}_{\mathrm{e}}^{\mathsf{H}}-\bar{U}_{\mathrm{e}}\hat{U}^{\mathsf{H}}-\hat{U}\bar{U}_{\mathrm{e}}^{\mathsf{H}}+\hat{U}\hat{U}^{\mathsf{H}}=(\bar{U}_{\mathrm{e}}-\hat{U})(\bar{U}_{\mathrm{e}}-\hat{U})^{\mathsf{H}}\succeq 0$, and hence, we have
@@
Inserting Inequality (75) in Inequality (E) yields $S_{\textnormal{exp-3}}(\epsilon,\tau_{3},U_{\mathrm{e}},\hat{U},\hat{W}_{\phi},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},D_{\mathrm{des}},\bar{D}_{3})\succeq 0$ (43c).